OpenStudy (anonymous):

simplify. 15+27+5+3+11+16+7+4

OpenStudy (anonymous):

plug it in a calculator and you should get the right answer

OpenStudy (anonymous):

do you know how to add two numbers together ?

OpenStudy (anonymous):

yah i just dont know what simplify means XD

OpenStudy (anonymous):

it means find the sum of all the numbers in this case

OpenStudy (anonymous):

15+5= 20 , 27+3=30 , 4+16=20 , 7+11= 18 so 20+30+20+18=88 hope that helped